OpenStudy (anonymous):

A real estate office handles 50 apartment units. When the rent is $720 per month, all units are occupied. However, on the average, for each $40 increase in rent, one unit becomes vacant. Each occupied unit requires an average of $48 per month for service and repairs. What rent should be charged to obtain a maximum profit?

OpenStudy (anonymous):

If n is the number of apartments are occupied. Total Rent company is getting: (720 + 40x(50-n)) x n - 48n We have to maximize the above equation.

OpenStudy (anonymous):

720n + 2000n - 40n^2 - 48n 2672n-40n^2 n= 33

OpenStudy (anonymous):

Rent = 720 + (50-33)x40 = 720 + 680 = $1400